OpenStudy (girlgoyleh.):

What is the value of w? 5(9 − 3) = w • 9 – 5 • 3

OpenStudy (jdoe0001):

http://www.mathgoodies.com/images/SS_PEMDAS_Goodies.jpg then solve for "w"

OpenStudy (girlgoyleh.):

i got 5 is that correct?

OpenStudy (jdoe0001):

hmm \(\bf 5(9 - 3) = w \cdot 9 - 5 \cdot 3\implies 5(4)=w9-15\implies 20=w9-15 \\ \quad \\ 20+15=w9\implies \cfrac{20+15}{9}=w\)

OpenStudy (jdoe0001):

hmmm wait.. shoot hold the mayo there

OpenStudy (girlgoyleh.):

im holding the mayo thanks

OpenStudy (jdoe0001):

and onions too, extra tomato and cheese =)

OpenStudy (jdoe0001):

\(\bf 5(9 - 3) = w \cdot 9 - 5 \cdot 3\implies 5(6)=w9-15\implies 30=w9-15 \\ \quad \\ 30+15=w9\implies \cfrac{\cancel{ 45 }}{\cancel{ 9 }}=w\implies 5=w\)

OpenStudy (girlgoyleh.):

but tomatoes and onions are the best part! thanks for your help:D

OpenStudy (jdoe0001):

heheh yw